OpenStudy (anonymous):

f(x) = 3x + 2; g(x) = 3x - 5 Find f/g

OpenStudy (ajprincess):

(f/g)(x)=f(x)/g(x)

OpenStudy (anonymous):

simple just divided them

OpenStudy (anonymous):

i dont know how ohugh i get -2/5 but thats not a choice

OpenStudy (anonymous):

should i show you the options

OpenStudy (ajprincess):

(f/g)(x)=f(x)/g(x) = (3x + 2)/ (3x - 5) Getting it @logynklode?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

but wouldnt that be -2/5

OpenStudy (anonymous):

no dont divide directly

OpenStudy (anonymous):

oh okay. so what do i do

OpenStudy (anonymous):

do they give you what is x? or just asking the function?

OpenStudy (ajprincess):

It wud be better if u post the options.

OpenStudy (anonymous):

if dont, then the answer coming from ajprincess

OpenStudy (anonymous):

thhat the whole question. no

OpenStudy (anonymous):

3x-5/3x+2 domain -2/3 3x+2/3x-5 domain 5/3 3x-5/3x+2 domain 5/3 3x+2/3x-5 domain -2/3

OpenStudy (anonymous):

so.

OpenStudy (anonymous):

what is domain in this mean?

OpenStudy (anonymous):

no idea. @matzen